# A Survey of Graph Unlearning

### v3

@@ Abstract
Keywords Graph Unlearning  $\cdot$ Graph Representation Learning  $\cdot$ Machine Unlearning  $\cdot$ Data Privacy  $\cdot$ Adversarial Attacks  $\cdot$ Recommender Systems  $\cdot$ Foundation Models
@@ 1 Introduction
In the ever-evolving landscape of artificial intelligence (AI) and data-driven decision-making, machine learning (ML) has emerged as a transformative force, empowering systems to learn from vast amounts of information and make intelligent predictions. However, as our digital world evolves, so do the ethical and privacy challenges it presents [1]. Enter the pioneering concept of machine unlearning—an awe-inspiring technological frontier designed to harmonize the power of ML with individual data rights and data privacy regulations [2, 3]. Machine unlearning represents an extraordinary leap forward in the pursuit of responsible AI, allowing us to gracefully erase the footprints of sensitive data from trained models, empowering individuals to exercise their right to be forgotten, and instilling confidence in AI’s ethical evolution [3].
An area where this ethical challenge becomes especially relevant is graph machine learning (GML). Unlearning plays a crucial role in GML applications, as GML is rapidly being adopted across various applications, including personalized recommendation systems and many others [4]. As an example of unlearning, consider a company that has trained a GML model on a social network data to suggest potential new friends based on user interactions and their features. If a user, such as Alice, requests her data be removed under privacy regulations, the model must undergo unlearning to eliminate the impact of her data [2]. Despite significant efforts to advance machine unlearning, the GML domain remains relatively unexplored in existing literature. This review, therefore, aims to address this gap by focusing on the latest developments under the umbrella of graph machine unlearning.
Graphs, are powerful mathematical structures, play a crucial role in representing complex systems [5]. From social networks, where nodes represent individuals and edges symbolize friendships, to biological networks like protein-protein interactions and transportation systems, graphs provide a compact and intuitive framework for data analysis [6]. Their versatility makes them ideal for modeling interconnected systems and extracting valuable insights for improved decision-making and forecasting [7, 8]. As such, understanding the graph structure is pivotal in tackling real-world problems.
Learning on graphs, or GML, represents a growing research area that focuses on leveraging graph-structured data to uncover patterns and make predictions [9, 10]. Unlike traditional ML methods, GML addresses the unique complexity and irregularity of graphs, excelling in tasks such as node classification, link prediction, and graph classification among others [11]. Furthermore, GML models have proven valuable in detecting communities, identifying anomalies, and offering promising solutions in fields like drug discovery, fraud detection, and social network analysis [12, 13]. By leveraging graph-based models, GML not only enhances our understanding of complex systems but also promises to revolutionize industries ranging from healthcare to smart city management [14, 15]. However, despite its growing importance, GML is still relatively new to industrial applications, and limited efforts have been made to explore machine unlearning in this context. As GML continues to evolve, unlearning in GML will likely become a crucial topic, given the broad applicability and privacy concerns associated with these systems.
Machine unlearning is a novel concept for ensuring right to be forgotten that refers to the process of selectively forgetting, or deleting specific data points or instances from the model [16]. Unlearning permits the exclusion of specific data from the training set of a model, making it more adaptable to evolving privacy regulations. Graph unlearning is an extension of machine unlearning designed specifically for graph-structured data [17]. It involves the complete removal of all traces of removed data such as node and edges from a graph, including all of its connections. In other terms, the objective of graph unlearning is to remove specific nodes and edges among others settings (see section 2) from the graph while preserving its overall structure. This level of data removal is significantly more complicated than traditional machine unlearning because it requires disentangling entities that are interconnected [18].
@@
The complexity of graph-structured data poses a challenge for unlearning in graphs. Relationships between nodes and edges are intricate in graphs, forming a dense web of interconnected information [19]. To unravel the impact of forgotten data while conserving the graph’s structure, sophisticated algorithms are required. These algorithms must be able to untangle these intricate connections without compromising the graph’s functionality [20]. In terms of data privacy and moral ML practices, graph unlearning is of the utmost relevance. As data privacy regulations and the rights of individuals to control their data become increasingly important, the ability to comprehensively remove sensitive information from graph-structured data becomes a crucial ethical safeguard. Ensuring the effective and complete eradication of forgotten data from graphs gives individuals more control over their information, instills confidence in ML systems, and upholds ethical standards for the responsible management of data [17].
@@
In the past few years, significant efforts have been made in the design of effective graph unlearning strategies, indicating a growing awareness of the significance of addressing data privacy and individual rights in the domain of GML [21, 22, 23]. These works have yielded optimistic advancements, demonstrating the researchers’ and practitioners’ commitment to tackling the complexities of graph unlearning. Despite these advancements, it is essential to recognize that the discipline is still relatively young and continues to evolve rapidly. As graph unlearning encounters the unique obstacles posed by interconnected networks, continued effort and creative solutions are required to overcome these obstacles. In light of the rapidly changing landscape of graph unlearning and the urgent need to resolve data privacy concerns in interconnected networks, we present this comprehensive survey as a crucial step in gaining a better understanding of the field. This survey seeks to summarize, analyze, and classify existing graph unlearning techniques, shedding light on their merits, limitations, and practical implications. Specifically, we collected papers by first gathering influential works and related surveys as seeds. From these, we expanded our survey by including more recent works that refer to those foundational papers. By collecting and analyzing these strategies, we hope to provide researchers and practitioners with a clear map for navigating this new domain. In addition, we envision this survey as a catalyst for future progress, outlining potential research and development avenues that will influence the future of graph unlearning. We offer the following contributions.
@@
  • We explore potential applications of graph unlearning across various fields and delve into intricate details for each case.
  • We provide a summary of existing open source models and data to help practitioners and inspired researchers.
  • Lastly, we emphasize multiple open research directions that demand further investigation, driving the pursuit of novel frontiers in this captivating domain.
@@ 2 Preliminaries
Generally, a graph is defined as $G=(V,E,X,X_{e})$, where $V$ represents a set of nodes denoted by $V=\{v_{1},v_{2},...,v_{n}\}$, and $E$ represents a set of edges denoted by $E=\{(v_{i},v_{j})\}$, signifying pairwise associations between nodes. The node feature matrix, $X\in\mathbb{R}^{|V|\times d}$, captures the attributes or characteristics of each node, where $d$ represents the dimensionality of node features. The edge feature matrix, $X_{e}\in\mathbb{R}^{|E|\times d^{\prime}}$, is optional and can be empty ($X_{e}=\emptyset$) in unweighted graphs. Here $d^{\prime}$ is the dimensionality of edge features [24]. Lets define $n$ the number of nodes and $m$ the number of edges in $G$. We also define $\mathcal{Y}=\{y_{1},y_{2},..,y_{n}\}$ consists of the labels associated with each node in the graph. The length of $|\mathcal{Y}|$ might be less than $n$ in some cases where fewer node labels are available. Consider the set of graphs or a graph dataset denoted as $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$ and its corresponding label set $L=\{l_{1},l_{2},\ldots,l_{N}\}$. Here $N$ denotes the number of samples (graphs) in the training data. Let $\mathcal{D}$ represent the corresponding embeddings of $\mathcal{G}$, given as $\mathcal{D}\in\mathbb{R}^{N\times h}$ and obtained through a graph embedding function denoted as $f(.)$. We introduce $\mathcal{D}^{\prime}$ as the embeddings obtained from $\mathcal{D}$ following a data removal request. Let’s also define $\mathcal{M}$, indicates the model trained on $\mathcal{D}$, $\mathcal{M}^{\prime}$ denotes the unlearned model updated through some unlearning method $A$ and $\hat{\mathcal{M}}$ indicates the model retrained from scratch after the removal request. We provide description of the all the notations used in this study in Table 1. In the following, we present a brief overview of each of these learning tasks:
@@
$X\in\mathbb{R}^{n\times d}$ The node feature matrix $X_{e}\in\mathbb{R}^{m\times d^{\prime}}$ Edge feature matrix
@@
$\mathcal{M}$ A model trained on $\mathcal{D}$ $\mathcal{M}^{\prime}$ An updated model for $\mathcal{D}^{\prime}$
$A$ A randomized mechanism for getting $\mathcal{M}^{\prime}$$\hat{\mathcal{M}}$ A model trained from scratch on $\mathcal{D}^{\prime}$
@@
[…]
$\mathcal{N}_{v}$ Node $v$’s neighbors $X_{v^{\prime}}$ The updated $X$ after $v$’s removal
$V^{-}$ A set of nodes requested for removal $H$ Hessian matrix
@@
Given a training graph $G$ and a set of labels $\mathcal{Y}$, node classification task aims to learn a representation vector $h_{v}$ for each node $v\in V$ using a function $f$ such that $v$’s label can be predicted as $y_{v}=f(h_{v})$, where $y_{v}\in\mathcal{Y}$
@@
In the node classification and link prediction tasks, it is common to work with a single graph $G$. The length of the label set $\mathcal{Y}$, in some cases, may be less than $n$ If the number of labeled nodes $|\mathcal{Y}|=n$, a fully supervised learning model can be trained. However, if $|\mathcal{Y}|<n$, a semi-supervised learning approach is often employed.
@@
Similarly, the graph regression and node regression problems share similarities with the graph classification and node classification tasks, but their primary focus lie in predicting continuous labels for each graph or node, rather than discrete classes. These machine learning tasks on graph-structured data, have garnered significant attention and found diverse applications in recent years. With these definitions in mind, we now turn our attention to defining the unlearning tasks on graphs.
Figure 1: Illustration of machine unlearning framework.
Definition 2.4 (Machine Unlearning)
Machine unlearning is a process that aims to update a trained machine learning model $\mathcal{M}$ in the presence of data removal, while mitigating the influence of the removed data point $\mathbf{x}_{i}$. Formally, let $\mathcal{D}$ be the original dataset with $N$ data points, where $\mathcal{D}=\{(\mathbf{x}_{1},y_{1}),(\mathbf{x}_{2},y_{2}),...,(\mathbf{x}_{N},y_{N})\}$, and $\mathcal{D}^{\prime}$ represents the dataset with the $i$-th data point removed. Machine unlearning aims to obtain an updated model $\mathcal{M}^{\prime}$ with minimal divergence from the original model $\mathcal{M}$ when trained on $\mathcal{D}^{\prime}$, while ensuring that the impact of $\mathbf{x}_{i}$ on $\mathcal{M}^{\prime}$ is significantly reduced.
We provide an illustration of machine unlearning in Figure 1. Initially, a learning method $\mathcal{M}$ is trained on a dataset $\mathcal{D}$. Once a removal request or a set of requests is received, the unlearning process is initiated. The unlearning method removes the impact of the requested data points and computes the updated model from $W$ to $W^{\prime}$, which corresponds to the parameters of the initially learned and unlearned models.
Machine unlearning can be categorized into two main categories: exact unlearning and approximate unlearning. In the following, we formally define these approaches.
Definition 2.5 (Exact Machine Unlearning)
Given a dataset $\mathcal{D}$ and a trained model $\mathcal{M}$, exact machine unlearning refers to the process of updating the model $\mathcal{M}$ after a removal request, such that the updated model $\mathcal{M}^{\prime}$ is identical to the model $\hat{\mathcal{M}}$, which is trained from scratch using the modified dataset $\mathcal{D}^{\prime}$. In other words, exact unlearning ensures that the retrained model $\mathcal{M}^{\prime}$ is indistinguishable from a model freshly trained on the altered dataset $\mathcal{D}^{\prime}$, effectively eliminating the influence of the removed data points.
Definition 2.6 (Approximate Machine Unlearning)
Given a dataset $\mathcal{\mathcal{D}}$ and a trained model $\mathcal{M}$, approximate machine unlearning refers to the process of updating the model $\mathcal{M}$ after a removal request in such a way that the updated model $\mathcal{M}^{\prime}$ is approximately equal to the model $\hat{\mathcal{M}}$, which is trained from scratch using the modified dataset $\mathcal{D}^{\prime}$. Unlike exact unlearning, where the retrained model is identical to the one trained from scratch on $\mathcal{D}^{\prime}$, approximate unlearning allows for small differences between the updated model $\mathcal{M}^{\prime}$ and the model $\hat{\mathcal{M}}$. The objective is to minimize the impact of the removed data points while accepting a certain level of discrepancy between the two models.
In approximate machine unlearning, the focus is on achieving a close resemblance between the retrained model and the one trained from scratch, rather than insisting on exact identity as in the case of exact unlearning. This flexibility allows for more practical scenarios where achieving perfect equality may be challenging or unnecessary, as long as the differences between the models remain within acceptable bounds. Now, let us proceed to define machine unlearning on graphs below.

### v4

@@ Abstract
{IEEEkeywords}
Adversarial Attacks, Data Privacy, Foundation Models, Graph Representation Learning, Graph Unlearning, Machine Unlearning, Recommender Systems
@@ 1 Introduction
In the ever-evolving landscape of artificial intelligence (AI) and data-driven decision-making, machine learning (ML) has emerged as a transformative force, empowering systems to learn from vast amounts of information and make intelligent predictions. However, as our digital world evolves, so do the ethical and privacy challenges it presents [1]. Enter the pioneering concept of machine unlearning—an awe-inspiring technological frontier designed to harmonize the power of ML with individual data rights and data privacy regulations [2, 3]. These policies allow us to gracefully erase the footprints of sensitive data from trained models, which has been proven to be easily recoverable through various adversarial attacks [4, 5, 6, 7]. Machine unlearning represents an extraordinary leap forward in the pursuit of responsible AI, empowering individuals to exercise their right to be forgotten [8, 9, 10], and instilling confidence in AI’s ethical evolution [11].
An area where this ethical challenge becomes especially relevant is graph machine learning (GML). Unlearning is an important technique in almost every application where GML models are utilized, including social networks, healthcare systems, financial networks, transportation networks, and more. As an example of unlearning, consider a company that has trained a GML model on a social network data to suggest potential new friends based on user interactions and their features. If a user requests their data be removed under privacy regulations, the model must undergo unlearning to eliminate the impact of their data [2]. Alternatively, from a security perspective, unlearning plays an essential role in addressing challenges and threats posed by adversarial attacks, which often involve malicious manipulations of nodes, edges, or features to compromise the functionality of GML models. In scenarios where adversaries inject harmful data into the graph structure–commonly referred to as poisoning attacks, it becomes crucial to effectively unlearn the malicious data without compromising benign portions of the graph or reducing the model’s predictive accuracy. Unlearning adversarial perturbations ensures the system’s integrity while safeguarding it from exploitation, maintaining its usability in critical applications such as recommendation systems, fraud detection, and network security [12].
@@
Though unlearning GML models have a great potential, the complexity of graph-structured data poses a challenge for unlearning in graphs. Relationships between nodes and edges are intricate in graphs, forming a dense web of interconnected information [13]. Similarly, GML uses message passing to aggregate information from neighboring nodes and edges. A single removed node, edge, or subgraph can influence the embeddings of distant nodes through multiple iterations of message passing. This makes it challenging to completely erase the influence of the removed data without retraining the model from scratch. Thus, to unravel the impact of forgotten data while conserving the graph’s structure, sophisticated algorithms are required. These algorithms must be able to untangle these intricate connections without compromising the graph’s functionality [14]. Moreover, as data privacy regulations and the rights of individuals to control their data become increasingly important, the ability to comprehensively remove sensitive information from graph-structured data becomes a crucial ethical safeguard. Ensuring the effective and complete eradication of forgotten data from graphs gives individuals more control over their information, instills confidence in ML systems, and upholds ethical standards for the responsible management of data [15].
@@
Significant efforts have been made in the design of effective graph unlearning strategies, indicating a growing awareness of the importance of addressing data privacy and individual rights in the domain of GML [16, 17]. These works have yielded notable advancements, demonstrating the researchers’ and practitioners’ commitment to tackling the complexities of graph unlearning. However, it is essential to recognize that the discipline is still relatively young and continues to evolve rapidly. As graph unlearning encounters the unique obstacles posed by interconnected networks, continued effort and creative solutions are required to overcome these obstacles.
In light of the rapidly changing landscape of graph unlearning and the urgent need to resolve data privacy concerns in interconnected networks, we present this comprehensive survey as a crucial step in gaining a better understanding of the field. This survey seeks to summarize, analyze, and classify existing graph unlearning techniques, shedding light on their merits, limitations, and practical implications. Specifically, we collected papers by first gathering influential works and related surveys as seeds. From these, we expanded our survey by including more recent works that refer to those foundational papers. By collecting and analyzing these strategies, we hope to provide researchers and practitioners with a clear map for navigating this new domain. In addition, we envision this survey as a catalyst for future progress, outlining potential research and development avenues that will influence the future of graph unlearning. We offer the following contributions.
@@
  • We explore potential applications of graph unlearning across various fields, and delve into intricate details of their current development.
  • We emphasize multiple open research directions that demand further investigation, driving the pursuit of novel frontiers in this captivating domain.
In recent years, a number of comprehensive survey papers have addressed the topic of machine unlearning [18, 19, 20, 19, 21]. Despite their valuable contributions, it is important to note that the aforementioned studies concentrate solely on general machine unlearning paradigms, which entails fundamentally different challenges and techniques than GML. Given that standard machine learning algorithms are not directly applicable to graph-structured data, unlearning methods designed for conventional settings cannot be straightforwardly transferred to graph learning scenarios. To the best of our knowledge, the only previous work that addressed graph unlearning is OpenGU [22]; however, this survey focuses mostly on benchmarking recent results and does not adequately represent the generate landscape of graph unlearning. This survey aims to bridge this gap by presenting a systematic and focused review of the emerging field of graph unlearning.
@@ 2 Preliminaries
Generally, a graph is defined as $G=(V,E,X,X_{e})$, where $V$ represents a set of nodes denoted by $V=\{v_{1},v_{2},...,v_{n}\}$, and $E$ represents a set of edges denoted by $E=\{(v_{i},v_{j})\}$, signifying pairwise associations between nodes. The node feature matrix, $X\in\mathbb{R}^{|V|\times d}$, captures the attributes or characteristics of each node, where $d$ represents the dimensionality of node features. The edge feature matrix, $X_{e}\in\mathbb{R}^{|E|\times d^{\prime}}$, is optional and can be empty ($X_{e}=\emptyset$) in unweighted graphs. Here $d^{\prime}$ is the dimensionality of edge features [23]. Lets define $n$ the number of nodes and $m$ the number of edges in $G$. We also define $\mathcal{Y}=\{y_{1},y_{2},..,y_{n}\}$ consists of the labels associated with each node in the graph. The length of $|\mathcal{Y}|$ might be less than $n$ in some cases where fewer node labels are available. Consider the set of graphs or a graph dataset denoted as $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$ and its corresponding label set $L=\{l_{1},l_{2},\ldots,l_{N}\}$. Here $N$ denotes the number of samples (graphs) in the training data. Let $\mathcal{D}$ represent the corresponding embeddings (dataset) of $\mathcal{G}$, given as $\mathcal{D}\in\mathbb{R}^{N\times h}$ and obtained through a graph embedding function denoted as $f(.)$.
Let us define $\theta^{o}=\mathcal{A}(\mathcal{D})$ as the model parameters obtained by applying the learning algorithm $\mathcal{A}$ to the full dataset $\mathcal{D}$. Let $\mathcal{S}\subseteq\mathcal{D}$ denote the forget set, and define $\mathcal{D}^{\prime}=\mathcal{D}\setminus\mathcal{S}$ as the dataset obtained after the removal request. We denote by $\theta^{r}=\mathcal{A}(\mathcal{D}\setminus\mathcal{S})$ the parameters of the retrained model obtained by applying $\mathcal{A}$ to the reduced dataset $\mathcal{D}^{\prime}$. Additionally, let $\theta^{u}=\mathcal{U}(\theta^{o},\mathcal{S},\mathcal{D})$ be the parameters of the unlearned model produced by an unlearning algorithm $\mathcal{U}$, which takes the original model $\theta^{o}$, the forget set $\mathcal{S}$, and the original dataset $\mathcal{D}$, and outputs $\theta^{u}$ via a post-processing mechanism.
2.1 General Machine Unlearning
Building on this notation, the primary objective of a machine unlearning method is to eliminate the influence of a specified forget set $\mathcal{S}$ from the trained model $\theta^{o}$.
Definition 2.1 $($$\varepsilon,\delta$)-unlearning [24])
For a fixed dataset $\mathcal{D}$, forget set $\mathcal{S}\subseteq\mathcal{D}$, and a randomized learning algorithm $\mathcal{A}$, an unlearning algorithm $\mathcal{U}$ is $(\varepsilon,\delta)$-unlearning with respect to $(\mathcal{D},\mathcal{S},\mathcal{A})$ if for all $R\subseteq\mathcal{R}$, where $\mathcal{R}$ denotes the output space (in this case, the space of model parameters $\theta$), we have:
$\Pr[\mathcal{A}(\mathcal{D}\setminus\mathcal{S})\in R]\leq e^{\varepsilon}\Pr[\mathcal{U}(\mathcal{A}(\mathcal{D}),\mathcal{S},\mathcal{D})\in R]+\delta,\text{and}$
$\Pr[\mathcal{U}(\mathcal{A}(\mathcal{D}),\mathcal{S},\mathcal{D})\in R]\leq e^{\varepsilon}\Pr[\mathcal{A}(\mathcal{D}\setminus\mathcal{S})\in R]+\delta.$
The above definition expresses the degree of success of an unlearning algorithm $\mathcal{U}$ as a function of a notion of divergence between the distributions of the two models $\theta^{r}$ and $\theta^{u}$. Initially, a learning method $\mathcal{\theta^{0}}$ is trained on a dataset $\mathcal{D}$. Once a removal request or a set of requests is received, the unlearning process is initiated. The unlearning method $\mathcal{U}$ removes the impact of the forget set $\mathcal{S}$ and computes the updated model $\theta^{r}$ or $\theta^{u}$ accordingly. Based on the above definition, machine unlearning methods can be broadly classified into two categories: exact unlearning and approximate unlearning. The unlearning is said to be exact when $\varepsilon=0$ and $\delta=0$, meaning the output distributions of the unlearning algorithm and retraining are identical. In contrast, non-zero values of $\varepsilon$ and $\delta$ characterize approximate unlearning, with the degree of approximation increasing as these parameters grow. Notably, when $\varepsilon$ and $\delta$ are sufficiently small, the distributions of the retrained model $\theta^{r}=\mathcal{A}(\mathcal{D}\setminus\mathcal{S})$ and the unlearned model $\theta^{u}=\mathcal{U}(\mathcal{A}(\mathcal{D}),\mathcal{S},\mathcal{D})$ become nearly indistinguishable:
$\lim_{\epsilon,\delta\rightarrow 0}Pr[\mathcal{U}(\mathcal{A}(\mathcal{D}),\mathcal{S},\mathcal{D})\in R]=\Pr[\mathcal{A}(\mathcal{D}\setminus\mathcal{S})\in R]\;\forall R\in\mathcal{R},$
indicating a high-quality unlearning outcome.
The reason behind the above classification is the degree to which the unlearned model replicates the distribution of a retrained model. In exact unlearning, the output of the unlearning algorithm is identically distributed to that of retraining from scratch on the pruned dataset, requiring $\varepsilon=0$ and $\delta=0$. This level of precision, while ideal, is often computationally infeasible in practice, e.g. large foundational models, and thus is usually reserved for privacy-critical applications. Approximate unlearning relaxes this constraint by allowing small, bounded differences between the two distributions, captured by non-zero $\varepsilon$ and $\delta$. This formulation enables practical implementations while still providing quantifiable control over residual influence.
@@ 2.1 General Machine Unlearning
$X\in\mathbb{R}^{n\times d}$ The node feature matrix $X_{e}\in\mathbb{R}^{m\times d^{\prime}}$ The edge feature matrix
@@
$\theta^{0}$ Weights of a model trained on $\mathcal{D}$ $\theta^{r},\theta^{u}$ Weights of the retrained/unlearned models for $\mathcal{D}^{\prime}$
$\mathcal{A}$ A learning algorithm $\mathcal{U}$ A machine unlearning method
@@
$\mathcal{N}_{v}$ Node $v$’s neighbors $X_{\setminus v}$ The updated $X$ after $v$’s removal
$V^{-},E^{-}$ A set of nodes/edges requested for removal $H$ Hessian matrix
2.2 Graph Machine Learning
We first introduce the relevant concepts on graph machine learning, which serves as background for the following sections.
@@ 2.2 Graph Machine Learning
Given a training graph $G$ and a set of labels $\mathcal{Y}$, the node classification task aims to learn a representation vector $h_{v}$ for each node $v\in V$ using a function $f$ such that $v$’s label can be predicted as $y_{v}=f(h_{v})$, where $y_{v}\in\mathcal{Y}$ .
@@
Similarly, graph regression and node regression problems share similarities with the graph classification and node classification tasks, but their primary focus lies in predicting continuous labels for each graph or node, rather than discrete classes. These machine learning tasks on graph-structured data have garnered significant attention and found diverse applications in recent years. With these definitions in mind, we now turn our attention to defining unlearning tasks on graphs.
2.3 Graph Unlearning